OpenStudy (anonymous):

If the cos(A) = 2/3 and the angle A is between 180 and 360 degrees, including 180 and 360 degrees, determine the sin(A) and the tan(A). Give a detailed explanation.

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

cosine is adjacent and hypotenues

OpenStudy (anonymous):

so right of the bat adjacent is 2 and hypotenuse is 3

OpenStudy (anonymous):

using pythagorean theorem, the third side is sqrt(5) ooo, ugly

OpenStudy (anonymous):

yes, it is

OpenStudy (anonymous):

now find the sine and tangent; sine=opposite/hypotenuse=sqrt(5)/3 tangent=opposite/adjacent=sqrt(5)/2

OpenStudy (anonymous):

give or take; I didn't really consider the degree constraints yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I am not good with geometry at all

OpenStudy (anonymous):

What would I do next?

OpenStudy (anonymous):

Smartone are you here?

OpenStudy (anonymous):

well, those are the numbers, and since the constraint is 180 to 360, I would make them all negative

OpenStudy (anonymous):

ok, thanks, if I attach a file, would you be able to see what is in it?

OpenStudy (anonymous):

Sin is negative between 180 and 360 degrees Tan is positive between 180 and 270 and negative between 270 and 360

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

I am confused about this problem.

OpenStudy (anonymous):

If cos = 2/3 then cosine is only positive between 270 and 360

OpenStudy (anonymous):

So make sin negative and tan negative with the answers above.

OpenStudy (anonymous):

but I do not have the answer, I do not know how to get it

OpenStudy (anonymous):

Are you all with me?

OpenStudy (anonymous):

darn guess noy

OpenStudy (anonymous):

not*